# Malware analysis assisted by AI with R2AI

### v1

@@ 1 Abstract
  • The cost is usually noticably lower than the salary of a malware analyst, but attention and guidance is needed to keep it under control in cases where the AI would naturally loop without showing progress.
@@ 2 Introduction
To assess the results of AI assistance, we compare the analysis reports of independant researchers on given IoT and Linux malware, with the analysis we get with AI assistance. Malware evolving rapidly, the samples we select for this study are all recent: mid 2024 to early 2025. This research is limited to IoT and Linux malware, but the results are expected to extend similarly to Windows and Mac OS malware. On Android, a few quick tests have shown AI assistance is less interesting because better decompilers exist. Malware for iOS, Flutter and other specific platforms would need to be investigated separately.
The paper is structured as followed. First, we discuss related art and how AI is used for malware analysis in other circumstances. Then, in section 4, we introduce r2 and its AI extension, r2ai. In section 5, we focus on dhow r2ai communicates with the AI. Finally, in section 6 and 7, we discuss costs and quality.
@@ 3 Related art
Radare2 https://www.radare.org is a set of command-line tools for reverse engineering and binary analysis. Its disassembler, r2, is scriptable, supports many different architectures and is widely used by researchers to inspect binaries.
@@ 7.1.2 Quality of analysis compared with a human only analysis
For several IoT and Linux malware of 2024 and 2025 (see Table 3), we compare the quality of analysis made by humans only, and with AI’s assistance. The “human-only” analysis is done by independant researchers [11] [12] [13]. The analysis assisted by AI was performed by us, using r2ai.
@@ 7.3.2 Time gain/time loss for analysis
In terms of speed, we tried to measure time gain or time loss induced by using AI. Measuring time to analyze binaries with precision is impossible: (1) we don’t know how much effort researchers spent [11] [12] [13], (2) our own AI-assisted analysis was slowed down by the very fact of writing conference articles, and (3) time to analyze a binary obviously depends on researcher’s skills, if s/he has already seen a similar sample, and the desired level of details.
@@ 9 Acknowledgments
We wish to thank Sergi Alvarez, radare2 maintainer for his implementation, reactivity and help. We also express our gratitude to Pr Ludovic Apvrille, for his detailed review of the paper, and several discussions on AI.

### v2

@@ 1 Abstract
  • The cost is usually noticeably lower than the salary of a malware analyst, but attention and guidance is needed to keep it under control in cases where the AI would naturally loop without showing progress.
@@ 2 Introduction
To assess the results of AI assistance, we compare the analysis reports of independent researchers on given IoT and Linux malware, with the analysis we get with AI assistance. Malware evolving rapidly, the samples we select for this study are all recent: mid 2024 to early 2025. While this research focuses specifically on IoT and Linux malware, the methodology and findings may be applicable to Windows or Mac OS malware analysis. However, when analyzing mobile platforms such as Android, iOS or cross-platform frameworks like Flutter, different approaches might be necessary, potentially yielding different results.
The paper is structured as followed. First, we discuss related art and how AI is used for malware analysis in other circumstances. Then, in section 4, we introduce r2 and its AI extension, r2ai. In section 5, we focus on how r2ai communicates with the language model providers. Finally, in section 6 and 7, we discuss costs and quality.
@@ 3 Related art
Radare2 https://www.radare.org is a set of command-line tools for reverse engineering and binary analysis. Its primary tool known as r2 is an interactive repl that can easily be scripted with different languages, natively supporting Javascript, and it supports many different architectures and binary file formats, in addition it is widely used by researchers to inspect binaries.
@@ 7.1.2 Quality of analysis compared with a human only analysis
For several IoT and Linux malware of 2024 and 2025 (see Table 3), we compare the quality of analysis made by humans only, and with AI’s assistance. The “human-only” analysis is done by independent researchers [11] [12] [13]. The analysis assisted by AI was performed by us, using r2ai.
@@ 7.3.2 Time gain/time loss for analysis
In terms of speed, we tried to measure time gain or time loss induced by using AI.
Measuring time to analyze binaries with precision is impossible: (1) we don’t know how much effort researchers spent [11] [12] [13], (2) our own AI-assisted analysis was slowed down by the very fact of writing conference articles, and (3) time to analyze a binary obviously depends on researcher’s skills, if s/he has already seen a similar sample, and the desired level of details.
@@ 9 Acknowledgments
We wish to thank Sergi Alvarez, radare2 author for his implementation, reactivity and help. We also express our gratitude to Pr Ludovic Apvrille, for his detailed review of the paper, and several discussions on AI.